OpenStudy (anonymous):

PLEASE Help, Completely lost! Find a polynomial f(x) of degree 3 with real coefficients and the following zeros. −3 , +2i

Miracrown (miracrown):

Let's start by looking at a simpler example... What if it said the polynomial has zeros of -3, 5 and -9? What would we do to get the polynomial?

OpenStudy (anonymous):

Since x is equal to -3 and +2i, Then itll start off as (x+3)(x-2i). Since there is an imaginary number you must need the opposite of it. So itll be (x+3)(x-2i)(x+2i)=f(x) Solve from there.

OpenStudy (anonymous):

Must start off with the imaginary numbers. (x-2i)(x+2i)(x+3)=(x^2-2i+2i-4i^2)(x+3)=(x^2+4)(x+3)=x^3+3x^2+3x^2+12=x^3+6x^2+12

OpenStudy (anonymous):

@Miracrown You just plug them in and factor them right? Resulting in (x+3)(x-5)(x+9)? Which = \[x ^{3}+7x ^{2}-33x-135\] Can't remember which step to take next...please keep in mind it's been a long day and I'm not at my sharpest right now haha @marie.serina09 I just tried to work it out on scratch paper and with some help from some videos, does that end up = \[x ^{3}+3x ^{2}+4x+12\] Is that my answer?

OpenStudy (anonymous):

Im pretty sure. I mightve missed a step cx

OpenStudy (anonymous):

Hold on..

OpenStudy (anonymous):

No worries!

OpenStudy (anonymous):

Yup youre answer is correct

OpenStudy (anonymous):

Woohoo! Awesome, thank you so much! Have a good night! :)

OpenStudy (anonymous):

Youre welcome and you too! :)